OpenStudy (anonymous):

Find the LCM (lowest common multiple) of 14y^3 and 28y^6

OpenStudy (anonymous):

\[14y^3 = 2\times7\times y \times y \times y\]and\[28y^6 = 2\times7\times2\times y \times y \times y \times y \times y \times y\] Notice that \[28y^6 = 2y^3(14y^3)\] Therefore, the LCM is \[28y^6\]